Project

General

Profile

Actions

Run Plan for 23rd June 2014

This run follows a request from ISIS to investigate discrepancies in the beam loss monitors.
It will not involve sending beam into the channel and will run as an expert shift.
The run will take place between 9am and 12pm.

Daily schedule

Month Friday 20th June 2014 Monday 23rd June 2014
estimated time Afternoon 8:30am - 12:30 pm
MOM P. Soler R. Bayes
Lead shifter Searcher1: Henry Nebrensky P. Hanlet
Shifter 2 Searcher2: Chris Heidt P. Smith
BLOC n/a P. Smith
Aim Lock up hall Test beam loss DAQ
Beam? No No
Additional Personnel ISIS Personnel : Bryan Jones, Sarah Fisher

Shifter hours of service:

  • Friday:
    • Search team searches/secures the Hall
    • All beam line magnets will be left on standby over the weekend
  • Weekend
    • MOM will visit MLCR twice each day of weekend (morning and evening) to monitor status of magnets and cooler/compressors in lieu of walk-through
    • If any issue arises, MOM will enter MICE Hall will be
  • Monday:
    • Shifter 1 - arrive 8:30 - finish 12:30pm
    • Shifter 2 - arrive 8:30 - finish 12:30pm

Data Requested

No data will be taken in this run

Beamline Configurations

We will run at 1V beam loss with the target in double dip mode.
The plan would be to dip the target at both MS/128 and MS/64 to ~ 1V level.
ISIS will then do BLM tests to find out what is wrong with the ISIS DAQ.
Sarah Fisher and Bryan Jones from ISIS will be present.

Henry: Confirm we have 2 ms headroom for increasing target delay at MS/64 without affecting following pulse - this is a <5min. job agreed verbally with Dean Adams (11 June 2014). Ideally do this at 4V MS/64 running - ISIS will need a couple of loss levels to understand the problem .

There will not be any running of beam into the hall, but we are still required to lock the Hall.

Run must finish at 12:00 to allow work on the DS power supply to start at 12:30.

Updated by Soler, Paul about 8 years ago ยท 10 revisions